What is Libermanns nitroso test ? Does only tert aromatic amines ans this test?

Grade:12

5 Answers

Aarti Gupta
9 years ago
Libermann’s nitroso test is is used as a test for secondary amines.Secondary amines(aliphatic as well as aromatic) reacts with nitrous acid to form N-nitrosoamines.
(CH3)2NH + HONO ---------------> (CH3)2N-N=O + H2O
Where (CH3)2 N-N=O is N-Nitrosodimethyl amine.Nitrosoamines are water soluble yellow oils and when warmed with phenol and few dropos of conc.H2SO4 produce a green colour solution which turns blue on adding alkali.This reaction is called Libermann’s nitroso reaction .Tertiary amine do not react with nitrous acid.
Pratishtha Sharma
4 years ago
The secondary amines show a distinguish test, known as Liebermann’s nitroso reaction, in which the N-nitrosoamine formed by rxn of secondary amine with HONO on treatment with phenol and conc.H2SO4 forms green clrd soln. which becomes blue clrd on addition of aq. NaOH and further on dilution turns red.
Yash Chourasiya
3 years ago
Dear Student

Libermann’s nitroso test is is used as a test for secondary amines.
Secondary amines(aliphatic as well as aromatic) reacts with nitrous acid to form N-nitrosoamines.
(CH3)2NH + HONO ---------------> (CH3)2N-N=O + H2O
Where (CH3)2 N-N=O is N-Nitrosodimethyl amine.
Nitrosoamines are water soluble yellow oils and when warmed with phenol and few dropos of conc.H2SO4 produce a green colour solution which turns blue on adding alkali.This reaction is called Libermann’s nitroso reaction .Tertiary amine do not react with nitrous acid.

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya
RISHIKA
3 years ago
DEAR FRIEND
Libermann's nitroso test is is used as a test for secondary amines. Secondary amines(aliphatic as well as aromatic) reacts with nitrous acid to form N-nitrosoamines. Where (CH3)2 N-N=O is N-Nitrosodimethyl amine. Nitrosoamines are water soluble yellow oils and when warmed with phenol and few dropos of conc.
THANK U
hamsika dornala
3 years ago
Liebermann's nitroso reaction is given by phenol,  secondary amine, and compounds containing nitroso  group